# KnowPhish: Large Language Models Meet Multimodal Knowledge Graphs for Enhancing Reference-Based Phishing Detection

### v1

@@ 8 Conclusions
In this work, we propose KnowPhish, a large-scale multimodal brand knowledge base covering more than 20k potential phishing targets, which can be integrated with any RBPD in a plug-and-play manner. We further propose KPD, a multimodal RBPD operating within both text and image modalities to detect phishing webpages with or without logos. Extensive experiments demonstrate the effectiveness and efficiency of KnowPhish and KPD under multiple settings. Moving forward, we foresee that integrating additional knowledge sources and LLM-related enhancements such as retrieval augmentation [30] can further enhance performance.
@@ Acknowledgement
Appendix A Additional Details on KnowPhish Construction
Phishing Targets of Different Industries
We provide more details of the brands from the eleven industries on each dataset in Table 8.
Industries
$D_{1}$
$D_{2}$
financial
Bank of America, Paypal, Credit Agricole, PostFinance
Bitkub, Credit Saison, Denizbank, Banco Do Brasil, GCash
online services
Outlook, Microsoft 365, Dropbox, Adobe, Onedrive
WeTransfer, Booking.com, Intuit, Biglobe, Mailchimp
telecommunication
AT&T, BT Group, Orange, Cox Communication
Shaw Communication, Swisscom, Singtel, Bell, Etisalat
e-commerce
Amazon, eBay, Rakuten, Americanas
Brooks Sports, Tesco, Loungefly, Shopee
social media
Instagram, Facebook, LinkedIn
WeChat, VKontakte
postal service
DHL, EMS, FedEx, La Poste
Australia Post, USPS, UPS, An Post, DPD
government
UK Gov, IRS, French Health Insurance,
Turkey Gov, Australia Gov, LTA Singapore
web portal
Google, Daum, AOL
Naver
video game
Steam, RuneScape, League of Legends
gambling
Bet365
other business
Delta Airline
KFC, AirNZ, Hydroqubec
Table 8: Examples of phishing targets belonging to different industries in $D_{1}$ and $D_{2}$
Wikidata Categories for KnowPhish Construction
We provide the full list of Narrow Categories $\mathcal{C}_{n}$ in Table 9 and General Categories $\mathcal{C}_{g}$ in Table 10. We put two Wikidata categories ‘online service’ and ‘government organization’ into $\mathcal{C}_{g}$ because we empirically find that it will lead to an excessively large number of brands. We handle this by conditioning on their popularity, which is identical to put them into $\mathcal{C}_{g}$.
Industries Wikidata Category Wikidata ID
financial
bank
Q22687
financial institution
Q650241
credit institution
Q730038
federal credit union
Q116763799
payment system
Q986008
digital wallet
Q1147226
cryptocurrency exchange
Q25401607
online service
webmail
Q327618
web service
Q193424
mobile app
Q620615
office suite
Q207170
telecommunication
telecommunication company
Q2401749
mobile network
Q15360302
mobile network operator
Q1941618
internet service provider
Q11371
e-commerce
online shop
Q4382945
online marketplace
Q3390477
social media
social media
Q202833
social networking service
Q3220391
online video platform
Q559856
postal service
postal service
Q1529128
package delivery
Q1447463
government
government
Q7188
web portal
web portal
Q186165
web search engine
Q4182287
video game
video game distribution platform
Q81989119
gambling
gambling
Q11416
Table 9: Full list of Narrow Categories $\mathcal{C}_{n}$
Industries Wikidata Category Wikidata ID
other business
business
Q4830453
public company
Q891723
enterprise
Q6881511
online service
Q19967801
government organization
Q2659904
Table 10: Full list of General Categories $\mathcal{C}_{g}$
KnowPhish Construction Algorithm
The complete KnowPhish construction algorithm is illustrated in Algorithm 2.
Input : Narrow Categories $\mathcal{C}_{n}$, General Categories $\mathcal{C}_{g}$, Wikidata Knowledge Graph $\mathcal{G}$, Top-ranked Domains $\mathcal{D}$, Max Domain Rank $\eta$
Output : Brand knowledge $\mathcal{B}$ with the name, logos, aliases, and domains of each brand
Notations : $r_{\mathcal{D}}(d)$ is the domain ranking of $d$ in $\mathcal{D}$, $h_{whois}(d)$ is the whois information for $d$, $\mathcal{N}(b)$ refers to the undirected neighbours of $b$ under ’owned by’ and ’parent organization’ relationship in $\mathcal{G}$
/* 1. Brand Search */
1 $\mathcal{B}_{n}\leftarrow\emptyset$, $\mathcal{B}_{g}\leftarrow\emptyset$;
2 for $c_{n}\in\mathcal{C}_{n}$ do
3       $\mathcal{C}_{n}^{\prime}\leftarrow\{c|(c,\texttt{subclass\_of},c_{n})\in\mathcal{G}\}$;
4       $\mathcal{B}_{n}(c_{n})\leftarrow\{b|(b,\texttt{instance\_of},c)\in\mathcal{G},c\in\{c_{n}\}\cup\mathcal{C}_{n}^{\prime}\}$;
5       $\mathcal{B}_{n}\leftarrow\mathcal{B}_{n}\cup\mathcal{B}_{n}(c_{n})$;
6
7 end for
8for $c_{g}\in\mathcal{C}_{g}$ do
9       $\mathcal{B}_{g}(c_{g})\leftarrow\{b|(b,\texttt{instance\_of},c_{g})\in\mathcal{G},\ r_{\mathcal{D}}(b.domains)\leq\eta\}$;
10       $\mathcal{B}_{g}\leftarrow\mathcal{B}_{g}\cup\mathcal{B}_{g}(c_{g})$;
11
12 end for
$\mathcal{B}$leftarrow $\mathcal{B}_{n}\cup\mathcal{B}_{g}$ /* 2. Knowledge Acquisition from Wikidata */
13 for $b\in\mathcal{B}$ do
14       $b.logos\leftarrow\{x|(b,\texttt{logo\_image},x)\in\mathcal{G}\}$;
15       $b.domains\leftarrow\{y.domain|(b,\texttt{official\_website},y)\in\mathcal{G}\}$;
16       $b.aliases\leftarrow\{z|(b,\texttt{label},z)\in\mathcal{G}\}$;
17
18 end for
/* 3. Knowledge Augmentation */
19 for $b\in\mathcal{B}$ do
       // Add logo variants
[…]
       // Add domain variants using Whois information
21       $b.domains\leftarrow b.domains\cup\{d|h_{whois}(d).org=h_{whois}(b.domains).org,\ d\in\mathcal{D}\}$;
22
23 end for
// Add domain variants via domain propagation
24 $\mathcal{B}^{\prime}\leftarrow\mathcal{B}$;
25 for $b^{\prime}\in\mathcal{B}^{\prime}$ do
26       $b^{\prime}.domains\leftarrow b^{\prime}.domains\cup\{b.domains|b\in\mathcal{N}(b^{\prime}),\ b\in\mathcal{B}\}$;
27
28 end for
$\mathcal{B}\leftarrow\mathcal{B}^{\prime}$;
30 return $\mathcal{B}$;
Algorithm 2 KnowPhish Construction
Appendix B Additional Details on KnowPhish Detector
Prompt Template for LLM Summary Generation Table 11 provides the complete prompt template to generate the LLM summary for the input webpage.
  Instruction: Define targeted brand as a brand that a webpage belongs to. Define credential-taking intention as a webpage’s intention to take users’ credentials, such as their email addresses, passwords, and so on. A credential-taking intention can be explicit or implicit, where explicit means having forms and input fields to submit user credentials directly, and implicit means not having explicit credential-taking intention, but instead having buttons or links redirecting users to another credential-taking webpage. Additionally, keywords related to user credentials, such as "Sign in", "Log in", "Register", "Account", "Assets", and "Password", are usually strong indicators of a credential-taking intention. Note that the texts in the HTML may be obfuscated into similar characters (e.g., ’a’ is obfuscated into ’$\alpha$’, or ’b’ is obfuscated into ’$\beta$’). If such obfuscation exists, please deobfuscate it and correctify your output. Given the URL and HTML of a webpage P, answer (1) What the targeted brand of P is. If it is not identifiable, put "Not identifiable". Extract the brand name only and do not include extra details such as affiliated products, countries, or additional abbreviations; (2) What forms or input fields to submit user credentials are present; (3) What buttons or links are present that redirect users to another credential-taking webpage; (4) What important keywords are present; (5) Whether there is a credential-taking intention; (6) Reason to the answer in (5). Start the answer to each of (1) to (6) on a new line.
URL: https://1staskyoude2-gopnumze9.top/
HTML: <title> Adobe-PDF Singapore sell everything you need </title> <a> </a> <a> </a> © 2023 Adobe. All brands are the property of their respective owners.
Answer:
(1) Adobe
(2) There are no forms or input fields to submit user credentials.
(3) There are no buttons or links directing the user to another credential-taking page.
(4) There are no important keywords.
(5) no intention
(6) The answer is according to (2), (3), and (4).
URL: https://cryptoinex.com/h5/
HTML: <title> Home - Cryptoin Online For Business - CPT </title> 本站点必须要开启JavaScript才能运行 Cryptoin currency Total assets equivalent (USD) 0.00 Announcement on Delisting SGB/USDT Token Pair Announcement on Delisting Selected Token Pairs Announcement on Supporting Ethereum London Hard Fork locked mining more 3day USDT lock up to earn coins 10 start 3% Daily rate of return 1day USDT lock up to earn coins 100 start 7% Daily rate of return 15day USDT lock up to earn coins 1000 start 15% Daily rate of return 30day USDT lock up to earn coins 10000 start 30% Daily rate of return 60day USDT lock up to earn coins 100000 start 60% Daily rate of return Quote more BTC/USDT -0.54% 22437.79000 DTC/USDT -0.49% 1570.48000 XRP/USDT +0.87% 0.37054 LTC/USDT -1.39% 87.52000 EOS/USDT -1.00% 1.21330 YMT/USDT -0.61% 1.26965 BCH/USDT +0.02% 124.50000 Quote more trading pair Latest Price Today change XRP /USDT 24H: 11681739.60651 0.37054 +0.87% LTC /USDT 24H: 5219.74600 87.52000 -1.39% EOS /USDT 24H: 204125.08287 1.21330 -1.00% YMT /USDT 24H: 490.15695 1.26965 -0.61% BCH /USDT 24H: 1079.22726 124.50000 +0.02% Home Markets Trade Finance Assets Cancel OK
Answer:
(1) Cryptoin
(2) There are no forms or input fields to submit user credentials.
(3) There are no buttons or links directing the user to another credential-taking page.
(4) There is a keyword "Assets" related to user assets.
(5) implicit intention
(6) The answer is according to (4).
URL: https://app.afe-n2jhk.com/index/login/login/token
HTML: English Chinese English <form> <inputtxt> </inputtxt> <inputpwd> </inputpwd> The account or password is wrong, please re-enter!! <button> Login </button> </form> <a> Open an account now </a> <a> Online service </a> <a> ©1998-2021 Rights Reserved </a> <a> AFE ©1998-2021 Rights Reserved </a>
Answer:
(1) AFE
(2) There is a form containing account and password input fields.
(3) There is a button with the label "Login" and a link with the text "Open an account now".
(4) There are keywords "account", "password", and "Login" related to user credentials.
(5) explicit intention
(6) The answer is according to (2).
URL: {U}
HTML: {H}
Answer:
Table 11: Prompt template to generate the LLM summary (including the text brand and CRP summary)
Estimated Cost for LLM Query
The LLM Summarizer in KPD leverages GPT-3.5-turbo-instruct as its LLM backbone. Here, we provide an estimation of the cost incurred by this LLM query. Take our TR-OP dataset as an example: on average, each webpage sample has 2,588 input tokens and 108 output tokens. With the API pricing at \$0.0015 per 1,000 input tokens and \$0.0020 per 1,000 output tokens, the estimated price for LLM summary per webpage sample is calculated as follows:
Cost $=\left(\frac{2588}{1000}\right)\times 0.0015+\left(\frac{108}{1000}\right)\times 0.002$
$\approx 0.0041\ \text{USD}$
Hence, the estimated cost for running KPD on the entire TR-OP dataset is approximately 41 USD.
Appendix C Additional Details on Experiments
Implementation Details
@@ Appendix C Additional Details on Experiments
Motivating Examples for HTML Obfuscation
Figure 12: A phishing webpage with its intended brand typosquatted.
Figure 12 provides an example of typosquatting, based on which we develop the HTML obfuscation techniques studies in this paper.
Logoless Phishing Webpages in Field Study
Figure 13 shows a few examples of logoless phishing webpages detected by KPD on SG-SCAN dataset.
Refer to caption Refer to caption
Figure 13: A few examples of logoless phishing webpages targeting WhatsApp and Singapore Post on SG-SCAN dataset.
Error Analysis
Figure 14 is an example of extremely implicit CRP that our text-based CRP classifier fails to detect.
Figure 14: A phishing webpage targeting Telegram with extremely implicit credential-requiring intention.

### v2

@@ 10 Conclusions
In this work, we propose KnowPhish, a large-scale multimodal brand knowledge base covering more than 20k potential phishing targets, which can be integrated with any RBPD in a plug-and-play manner. We further propose KPD, a multimodal RBPD operating within both text and image modalities to detect phishing webpages with or without logos. Extensive experiments demonstrate the effectiveness of KnowPhish and KPD, and highlight the deployment efficiency of KnowPhish over DynaPhish across multiple settings. Moving forward, we foresee that integrating additional knowledge sources and LLM-related enhancements such as retrieval augmentation [31] can further enhance performance.
@@ Acknowledgement
Appendix A Appendix
A.1 Implementation Details
@@
A.2 Logo-less Phishing in Field Study
Figure 11 shows a few examples of logo-less phishing webpages detected by KPD+KnowPhish on SG-SCAN dataset.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: A few examples of logo-less phishing webpages targeting WhatsApp, Singapore Post, Australia Post, and Amazon on SG-SCAN dataset.
A.3 False Positives Analysis in Field Study
Here, we discuss two primary types of false positives that arise from DynaPhish: the inclusion of web-hosting brands and the popularity validation failure of benign domains. They account for more than 70% of the false positives. We elaborate on each:
Inclusion of web-hosting brands
Figure 12 shows a few false positive examples resulting from the inclusion of web-hosting brands, such as file-hosting brands Nextcloud, FileGator, and a domain hosting brand Bitly. These benign webpages display the logos of web-hosting brands simply because they utilize their services, not because they are conducting phishing attacks. However, the RBPDs mistakenly report them as phishing due to the inconsistency between the logos and their legitimate domains. This issue also exists in the original BKB of Phishpedia/PhishIntention, which includes a domain hosting brand GoDaddy.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Examples of false positives by DynaPhish and Phishpedia/PhishIntention’s original BKB on SG-SCAN dataset, arising due to web-hosting brands.
A straightforward solution is to exclude these brands from the BKB. In our work, KnowPhish adopts a postprocess operation that filters these brands out by conditioning on their Wikidata categories (i.e., any brand that belongs to ‘file synchronization’, ‘URL shortener’, ‘blog’, or ‘domain name registrar’ is excluded from the brand list $\mathcal{B}$). However, since the Wikidata information may be incomplete, we foresee that a more comprehensive list of such web-hosting brands should be collected for exclusion.
Popularity validation failure of benign domains
DynaPhish also experiences false positives when failing to validate the popularity of benign domains. For example, googleadservices.com and documentforce.com are the legitimate domains of Google and Salesforce, respectively. However, popularity validation fails when DynaPhish Google Searches with these domains as queries, because they are not included in the search results. Since the popularity validation fails, and DynaPhish identifies the brand intention from the logos, phishing is mistakenly reported.
A.4 Potential Defense Improvements against Adversarial Attacks
As discussed in Section 6.5, our defense partially mitigates the adverse effect of prompt injection and text-to-image attacks, and we believe future studies are needed to better address these problems. Potential improvements for prompt injection defense include handling the adversarial prompts at input stage[23], inference stage[6], or both[36]. Regarding text-to-image attacks, additional components such as OCR with well-crafted prompts can assist LLMs to better analyze the texts from screenshots, as supervised OCR models have been found to outperform our defense model GPT-4V in OCR tasks[52].